OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1=2, an=-3an-1

OpenStudy (anonymous):

\[a_2=-3\times a_{1}=-3\times 2=-6\]

OpenStudy (anonymous):

similarly \[a_3=-3\times a_2=-3\times -6=18\] and so on

OpenStudy (anonymous):

Thank you for always helping me! I really appreciate it! :)

OpenStudy (anonymous):

I will be posting more so yea. If you could help with those too, I would really appreciate it! :)